### v2

# Solving systems of equations on antichains for the computation of the ninth Dedekind Number

Abstract
We study four systems of equations, together with a way to count the number of solutions. One of the results was used in the recent computation of D(9), the others have potential to speed up existing techniques in the future.
keywords:
Dedekind numbers, antichains, algebra of distributive lattices, posets, intervals
journal: Journal of  Templates
1 Introduction
Calculating Dedekind Numbers (OEIS series A000372 [1]) has always followed progress in both the Mathemathical domain, and the Computational Implementation Domain. Computing Subsequent Numbers is known as ”Dedekind’s Problem”. Since the year 1991 and up until March 2023, the highest known term in this sequence was D(8), computed by Doug Wiedemann [2] on a Cray 2 supercomputer. In April 2023 two independent computations of D(9) were announced (see [3, 4]). The computations happened completely independently, were based on different methodologies and relied on very different hardware. Christian Jäkel used a GPU infrastructure and built on a matrix formulation, which allowed to use the advanced techniques and highly developed software for numerical analysis. In our computation (Van Hirtum et al. [4]), we, together with our co-authors, relied on the advanced super computing infrastructure at the Pc2 centre at Paderborn University, using cutting edge FPGA hardware and an ingenious implementation of the so-called P-coefficient formula [5]. At the end of the latter computation, some uncertainty due to the semi-experimental setting of the FPGA infrastructure was identified. Although we obtained the result on March 8, it was only when the result of Jäkel came out, confirming our result, that we could be certain. An extensive report on this computational process is in [4], concentrating on the very important computational aspects and the ingenious implementation which brought important new insights for FPGA technology. The present paper elaborates on the mathematical aspects, with a discussion on the P-coefficient and possible extensions.
2 Definitions and notations
2.1 The Dedekind number
For any finite positive number $n\geq 0$, the n-th Dedekind number counts the number of monotone Boolean functions, or equivalently, the number of antichains, on the subsets of a finite set $\mathcal{A}=\{1,\dots,n\}$. To derive the properties in this paper, we will use the concept an antichain over $\mathcal{A}$, i.e. sets of pairwise non-inclusive subsets of $\mathcal{A}$.
The set of permutations of the elements of base set $A$ generates an equivalence relation on $\mathcal{D}_{n}$. The set of equivalence classes of this relation is denoted by $\mathcal{R}_{n}$ and the number of such equivalence classes is denoted by $R(n)$.
The known Dedekind numbers are shown in Table 1, including the result of our computation. Table 2 shows the known numbers $R(n)$ of equivalence classes of monotone Boolean functions under permutation of the elements of the base set. Note that the last result dates from 2023.
D(0) 2 Dedekind (1897)
D(1) 3 Dedekind (1897)
D(2) 6 Dedekind (1897)
D(3) 20 Dedekind (1897)
D(4) 168 Dedekind (1897)
D(5) 7581 Church (1940)
D(6) 7828354 Ward (1946)
D(7) 2414682040998 Church (1965)
D(8) 56130437228687557907788 Wiedemann (1991)
D(9) 286386577668298411128469151667598498812366 Jäkel, Van Hirtum et al. (2023)
Table 1: Known Dedekind Numbers, OEIS series A000372 [1], including our result.
R(0) 2
R(1) 3
R(2) 5
R(3) 10
R(4) 30
R(5) 210
R(6) 16353
R(7) 490013148 Tamon Stephen & Timothy Yusun (2014) [6]
R(8) 1392195548889993358 Bartelomiej Pawelski (2021) [7]
R(9) 789204635842035040527740846300252680 Bartelomiej Pawelski (2023)
Table 2: Known Equivalence Class Counts, OEIS series A003182 [8]
2.2 Notation
Let $n\in\mathbb{N},A=\{1,\dots,n\},2^{A}=\{X\subseteq A\},\mathbb{B}=\{false,true\}$.
  • Definition
    A Boolean function $f:2^{A}\rightarrow\mathbb{B}$ is monotone iff
    $\forall X,Y\in 2^{A}:X\subseteq Y\Rightarrow f(Y)\Rightarrow f(X)$
Note that a monotone Boolean function is completely defined by the set of sets which are maximal among the sets for which the function value is true.
  • Definition
    A set $X\in 2^{A}$ is a maximal set of a monotone Boolean function $f$ iff
    $f(X)=true\ and\ \forall Y\in 2^{A}:X\subsetneq Y\Rightarrow f(Y)=false$
For any monotone Boolean function, no two of its maximal sets include each other. A set of sets with this property is called an antichain:
  • Definition
    A set of sets $\sigma\subseteq 2^{A}$ is an antichain iff
    $\forall X,Y\in\sigma:X\not\subseteq Y\ and\ Y\not\subseteq X$
In what follows, we will represent antichains by letters from the Greek alphabet $\alpha,\beta,\ldots$. The elements of the set $A$ will be denoted by Latin minuscules $a,b,c,\ldots$ or by digits $1\ldots 9$. For subsets, unless there is a possibility for confusion, we will use the abbreviation $abc\equiv\{a,b,c\},123\equiv\{1,2,3\},0\equiv\bot$ or refer to them by by Latin majuscules $X,Y,\ldots$.
  • Definition
    An antichain $\alpha$ is less or equal than an antichain $\beta$ iff each set in $\alpha$ is included in at least one set of $\beta$:
    $\alpha\leq\beta\Leftrightarrow\forall X\in\alpha:\exists Y\in\beta:X\subseteq Y$
  • Notation
    We denote by $\bot$ and $\top$ the smallest, respectively the largest, element of $D_{n}$:
    $\forall\alpha\in D_{n}:\bot\leq\alpha\leq\top$ (1)
  • Definition
    For any two elements $\alpha,\beta\in\mathcal{D}_{n}$, the interval with bottom $\alpha$ and top $\beta$ is the set of antichains $\chi$ satisfying $\alpha\leq\chi\leq\beta$:
    $\forall\alpha,\beta\in\mathcal{D}_{n}:[\alpha,\beta]=\{\chi\in\mathcal{D}_{n}:\alpha\leq\chi\leq\beta\}$
  • Definition
    For $\alpha,\beta\in\mathcal{D}_{n}$, the join $\alpha\vee\beta$ and the meet $\alpha\wedge\beta$ are the antichains defined by
    $\alpha\vee\beta=max(\{X\in\alpha\cup\beta\})$ (2)
    $\alpha\wedge\beta=max(\{X\cap Y|X\in\alpha,Y\in\beta\})$ (3)
Where $max$ denotes the maximum w.r.t. inclusion of a set of subsets of $\mathcal{A}$, i.e.
$S\subset 2^{A}\Rightarrow max(S)=\{X\in S:\forall Y\in S:X\not\subsetneq Y\}$
Finally, in the formulas below, a number defined for each pair $\alpha\leq\beta\in\mathcal{D}_{n}$ plays an important role. We refer to this number as the connector number $C_{\alpha,\beta}$. It counts the number of connected components of a graph with sets in $\beta$ as vertices and edges between two sets if the intersection is not dominated by $\alpha$ (exact definition follows).
3 Related Work
The work in this article builds upon three papers by Patrick De Causmaecker et al. [9] [5], [10]. Recently, a number of interesting results were published by Bartolomiej Pawelski, e.g. the recent computation of $R(9)$ [11].
Finally of course we want to mention the work of Christian Jäkel, who computed the 9th Dedekind Number at about the same time as us, but using different formula’s and different hardware. [3].
Central to our computation was a result on the number of solutions of the set of two equations
$\chi\vee\upsilon=\beta$ (4)
$\chi\wedge\upsilon=\alpha$ (5)
for $\chi,\upsilon\in D_{n}$ is given by $2^{C_{\alpha,\beta}}$. This is called the P-Coefficient also denoted as PCoeff [5, 10]. Below we will elaborate on this result and study its extension to larger systems of equations.
In the following sections, we derive various P-coefficient formulas.
4 A system of two equations
Let $\alpha\leq\beta$ be two antichains. Consider the set of equations
$\chi\wedge\upsilon=\alpha$ (6)
$\chi\vee\upsilon=\beta$ (7)
It is clear that the system has no solution if $\alpha\not\leq\beta$.
Lemma 4.1
In any solution of System 4, $\chi=\alpha\vee\chi_{1},\upsilon=\alpha\vee\upsilon_{1}$ with $\chi_{1}\subseteq\beta-\alpha,\upsilon_{1}\subseteq\beta-\alpha$.
  • Proof
    Suppose $X\in\chi_{1},X\not\in\beta$. Due to equation 7, there must be a set $X^{\prime}\in\upsilon$ such that $X\subseteq X^{\prime}$ and $X^{\prime}\in\beta$. Due to equation 6, this implies $X\in\alpha$ in contradiction with $\chi_{1}\subseteq\beta-\alpha$. Similar arguments work for $\upsilon_{1}$. $\qed$
Lemma 4.2
For any set $A\in\beta-\alpha$, we have that $A\in\chi-\upsilon$ or $A\in\upsilon-\chi$.
  • Proof
    From equation 7, it follows that $A$ must be in at least one of $\chi$ or $\upsilon$. From the same equation it follows that no set dominating $A$ can be in any of the two. If $A$ were in both $\chi$ and $\upsilon$, $A\in\chi\wedge\upsilon$ and equation 6 would be violated. $\qed$
Lemma 4.3
Let $A,B\in\beta,A\not\in\alpha,B\not\in\alpha$, such that $\{A\cap B\}\not\leq\alpha$. We have
$A\in\chi\Leftrightarrow B\in\chi$
$A\in\upsilon\Leftrightarrow B\in\upsilon$
  • Proof
    Due to Lemma 4.2, $A$ and $B$ are each in exactly one of $\chi$ and $\upsilon$. If they were each in a different one, we would have $\{A\cap B\}\leq\chi\wedge\upsilon=\alpha$ due to equation 6. A contradiction. $\qed$
  • Definition
    Given two antichains $\alpha\leq\beta$, two sets $A,B\in\beta-\alpha$ are said to be directly connected with respect to $\alpha$ iff $\{A\cap B\}\not\leq\alpha$.
  • Definition
    Two sets $A,B\in\beta-\alpha$ are said to be connected w.r.t. $\alpha$ iff $A$ and $B$ are directly connected w.r.t. $\alpha$ or there exist $Z_{1},Z_{2},\ldots,Z_{x}\in\beta$, $x\geq 1$, such that $A$ is directly connected with $Z_{1}$, $B$ is directly connected with $Z_{x}$ and each $Z_{i}$ is directly connected to $Z_{i+1}$ for $i\in\{1,\ldots,x-1\}$, all w.r.t. $\alpha$.
  • Definition
    For two antichains $\alpha\leq\beta$, the connection graph $C_{\alpha}(\beta)$ is the graph with the sets of $\beta$ as vertices and edges
    $E=\{(A,B)|A,B\in\beta,A,B\text{ are connected w.r.t. }\alpha\}$
Theorem 4.4
Let $\alpha\leq\beta$ be two antichains, $(\chi_{0},\upsilon_{0})$ a partition of $\beta-\alpha$ such that any two sets $A,B\in\beta-\alpha$ that are connected w.r.t. $\alpha$ are in the same component of the partition. With the partition $(\chi_{0},\upsilon_{0})$ corresponds exactly one solution of System 4 and any solution of System 4 corresponds with exactly one such partition.
  • Proof
    Due to Lemma 4.1, any solution is determined by two subsets of $\beta-\alpha$. Due to lemma 4.3, any two directly connected sets $A,B\in\beta$ must come in the same antichain $\chi$ or $\upsilon$. By transitive closure, this must be true for any two connected sets $A,B\in\beta$ and hence for all sets in any connected component of $C_{\alpha}(\beta)$. Given a partition $(\chi_{0},\upsilon_{0})$ of $\beta-\alpha$ satisfying the criteria of the theorem, we arrive at a solution $\chi=\chi_{0}\vee\alpha,\upsilon=\upsilon_{0}\vee\alpha$. The reverse follows by identifying $\chi_{0}=\chi-\alpha,\upsilon_{0}=\upsilon-\alpha$. $\qed$
Corollary 4.5
The number of solutions of System 4 is given by
$2^{C_{\alpha,\beta}}$
where $C_{\alpha,\beta}$ is the number of connected components of the connection graph $C_{\alpha}(\beta)$.
  • Proof
    The number of partitions in two sets of the connected components of $C_{\alpha}(\beta)$ is given by $2^{C_{\alpha,\beta}}$. As was proven in Theorem 4.4, this is the number of solutions of System 4.
Corollary 4.6
The number of solutions to system 4 can be computed in polynomial time on the number of sets in $\beta-\alpha$.
  • Proof
    By Theorem 4.4, the problem is reduced to counting the connected components in an undirected graph with sets in $\beta-\alpha$ as vertices. This can be done in $O(V+E)$ time where $V$ is the number of vertices and $E$ is the number of edges. $\qed$
¡ltx:note¿T¡/ltx:note¿he number $2^{C_{\alpha,\beta}}$ is referred to as the P-coefficient in [4].
5 Application of System 4
System 4 can be used to compute the $(n+2)^{th}$ Dedekind number ($D(n+2)$) based on intervals in the space of antichains over $n$ elements.
Theorem 5.1
Let $n\geq 0$ be a natural number, $N=\{1,..,N\}$, $\mathcal{A}_{n}$ the lattice of antichains over $N$. We have
$D(n+2)=\sum_{\alpha\leq\beta\in\mathcal{D}_{n}}2^{C_{\alpha,\beta}}\times|[\bot,\alpha]||[\beta,\top]|$
  • Proof
    We use the decomposition for $\eta\in\mathcal{D}_{n+2}$ in terms of elements $\eta_{X}$ of $\mathcal{D}_{n}$:
    $\eta=$ $\eta_{0}\times\{12\}$
    $\vee\eta_{1}\times\{1\}\vee\eta_{2}\times\{2\}$
    $\vee\eta_{12}\times\{0\}$
    were sets such as $\{1,2\}$ are denoted by $12$ and $\bot$ by 00 as usual. The decomposition is unique iff $A\subseteq B\Rightarrow\eta_{A}\leq\eta_{B}$. We identify $\eta_{1},\eta_{2}$ with the variables $\chi,\upsilon$ in System 4. Any solution of System 4 for specific values of $\alpha,\beta$ will now correspond to exactly one antichain $\eta$ for which $\eta_{0}\in[\bot,\alpha]$, $\eta_{12}\in[\beta,\top]$. $\qed$
Corollary 5.2
$D(n+2)$ can be computed as a sum of $D(n-1)$ terms, each of which is the product of the sizes of two intervals in $\mathcal{D}_{n}$.
  • Proof
    The condition $\alpha\leq\beta$ for $\alpha,\beta\in\mathcal{D}_{n}$ allows exactly $D(n-1)$ terms.$\qed$
This improves on the number of terms used in the computation of $D(8)$ by Wiedemann [2], because without the condition $\alpha\leq\beta$, the number of terms is $D(n-1)^{2}$. Theorem 5.1 and its corollary 5.2 were the basis for our computation of $D(9)$ [4]. Of course, as Wiedemann did as well, we had to reduce the number of terms further using permutation symmetry.
6 A system of four equations
Let $\alpha\leq\beta_{1},\beta_{2},\beta_{3}$ be four antichains with $\alpha\leq\beta_{i}$ for $i\in{1,2,3}$. Consider the set of equations
$\chi_{1}\wedge\chi_{2}\wedge\chi_{3}=\alpha$ (8)
$\chi_{2}\vee\chi_{3}=\beta_{1}$ (9)
$\chi_{1}\vee\chi_{3}=\beta_{2}$ (10)
$\chi_{1}\vee\chi_{2}=\beta_{3}$ (11)
  • Notation
    We will use $\beta=\beta_{1}\cup\beta_{2}\cup\beta_{3}$.
Lemma 6.1
In any solution of System 8, $\chi_{i}=\alpha\vee\upsilon_{i}$ with $\upsilon_{i}\subseteq\beta-\alpha$.
  • Proof
    Suppose $X\in\chi_{1},X\not\in\beta,X\not\in\alpha$. Due to equation 10, there must be a set $X^{\prime}\in\chi_{3}$ such that $X\subseteq X^{\prime}$ and $X^{\prime}\in\beta_{2}$ and due to equation 11, there must be a set $X^{\prime\prime}\in\chi_{2}$ such that $X\subseteq X^{\prime\prime}$ and $X^{\prime\prime}\in\beta_{3}$. Due to equation 8, this implies $X\in\alpha$ in contradiction with $X\not\in\alpha$. Similar arguments work for $\chi_{2}$ and $\chi_{3}$. $\qed$
Lemma 6.2
For System 8 to have a solution, each set in $\beta-\alpha$ must be dominated by at least 2 of $\beta_{1},\beta_{2},\beta_{3}$.
  • Proof
    Let $A\in\beta_{1}-\alpha$, so that $\{A\}\leq\beta_{1}$ From equation 9, it follows that $A\in\chi_{2}$ or $A\in\chi_{3}$. In the first case, due to equation 11, we have $\{A\}\leq\beta_{3}$. In the second case, due to equation 10, we have $\{A\}\leq\beta_{2}$. $\qed$
Lemma 6.3
Any set in $\beta-\alpha$ dominated by exactly 2 of $\beta_{1},\beta_{2},\beta_{3}$ occurs in exactly one of $\chi_{1},\chi_{2},\chi_{3}$ for any solution of System 8, and there is only one possibility.
  • Proof
    Let $A\in\beta_{1}-\alpha$ be dominated by $\beta_{2}$ and not by $\beta_{3}$. Due to equation 11, $A$ cannot be dominated by $\chi_{1}$ or $\chi_{2}$. Due to equations 9 and 10, $A$ must be in $\chi_{3}$. $\qed$
Corollary 6.4
Any set $A\in\beta-\alpha$ dominated by exactly 2 of $\beta_{1},\beta_{2},\beta_{3}$ must occur in exactly 2 of $\beta_{1},\beta_{2},\beta_{3}$ for System 8 to have a solution.
  • Proof
    Let $A\in\beta_{1}-\alpha$ be dominated by $\beta_{2}$ and not by $\beta_{3}$. From the proof of Lemma 6.3, it follows that $A\in\chi_{3}$ cannot be dominated by $\chi_{1}$ or $\chi_{2}$. Due to equations 9 and 10, $A$ must be in $\beta_{2}$. $\qed$
Lemma 6.5
Let $A\in\beta-\alpha$, $A\in\beta_{1}\cap\beta_{2}\cap\beta_{3}$. $A$ must occur in exactly two of $\chi_{1},\chi_{2},\chi_{3}$.
  • Proof
    Equations 9, 10 and 11 imply that $A$ must be in at least two of $\chi_{1},\chi_{2},\chi_{3}$. Equation 8 implies that it cannot be in all three of $\chi_{1},\chi_{2},\chi_{3}$. $\qed$
Lemma 6.6
Let $A\in\beta-\alpha$, $A\leq\beta_{1},\beta_{2},\beta_{3}$. $A$ must be dominated by exactly two of $\chi_{1},\chi_{2},\chi_{3}$.
  • Proof
    Equations 9, 10 and 11 imply that $A$ must be dominated by at least two of $\chi_{1},\chi_{2},\chi_{3}$. Equation 8 implies that it cannot be dominated by all three of $\chi_{1},\chi_{2},\chi_{3}$. $\qed$
Lemma 6.7
Let $A,B\in\beta-\alpha$ both be dominated by $\beta_{1},\beta_{2},\beta_{3}$ and suppose $\{A\cap B\}\not\leq\alpha$. In any solution of System 8, exactly two of $\chi_{1},\chi_{2},\chi_{3}$ dominate both $A$ and $B$ while the third one does not dominate any of the two.
  • Proof
    From Lemma 6.6 it follows that $A$ as well as $B$ are dominated by exactly two of $\chi_{1},\chi_{2},\chi_{3}$. Suppose $A$ is dominated by $\chi_{1},\chi_{2}$ and $B$ is dominated by $\chi_{1},\chi_{3}$. Then $\{A\cap B\}\leq\chi_{1}\wedge\chi_{2}\wedge\chi_{3}=\alpha$ in contradiction with $\{A\cap B\}\not\leq\alpha$. $\qed$
  • Notation
    We denote by
    $\mathcal{A}=\{A\in\beta-\alpha|\{A\}\leq\beta_{1}\wedge\beta_{2}\wedge\beta_{3}\}$
    the set of sets that are dominated by all $\beta_{i}$ for $i\in\{1,2,3\}$.
  • Notation
    We denote by $\mathcal{C}_{i}$ the set of sets which must only be in $\chi_{i}$ for $i\in\{1,2,3\}$ in any solution of System 8 according to Lemma 6.3.
  • Definition
    For any connected component $g$ of the graph $C_{\alpha}(\mathcal{A})$, the weight of $g$ is defined as
    $w(g)=3-|\{i\in\{1,2,3\}|\exists C\in\mathcal{C}_{i}:\exists G\in g:\{C\cap G\}\not\leq\alpha\}|$
Theorem 6.8
Given $\alpha,\beta_{1},\beta_{2},\beta_{3}$ with $\alpha\leq\beta_{i}$ for $i\in\{1,2,3\}$ and $\beta=\beta_{1}\cup\beta_{2}\cup\beta_{3}$, the number of solutions of System 8 is given by
$P_{3}(\alpha,\beta_{1},\beta_{2},\beta_{3})=\prod_{\begin{subarray}{c}g\text{ connected component}\\ \text{ of }C_{\alpha}(\mathcal{A})\end{subarray}}\ w(g)$
  • Proof
    Due to Lemma 6.1, for all i, $\chi_{i}=\alpha\vee\upsilon_{i},\upsilon_{i}\subseteq\beta$ so that we only need to count subsets of $\beta$ leading to a solution of System 8. Let $g$ be a connected component of $C_{\alpha}(\mathcal{A})$. From Lemma 6.7, we can conclude that in any solution of System 8 there must be exactly two $i,j\in\{1,2,3\}$ such that $\chi_{i}$ and $\chi_{j}$ dominate all sets in $g$. The values for $i$ and $j$ can be chosen independently for all connected components, so there are potentially $3$ possibilities for any $g$. If there is a set $G\in g$ and a set $C\in\mathcal{C}_{i}$ such that $\{G\cap C\}\not\leq\alpha$, then all sets in $g$ must be dominated by $\chi_{i}$ reducing the number of possibilities by $1$. The number of values $i\in\{1,2,3\}$ for which this is true must be subtracted from $3$ to obtain the number of possibilities for $g$, resulting in the weight $w(g)$ defined above. $\qed$
Corollary 6.9
The number of solutions of System 8 can be computed in polynomial time in the number of sets in $\beta-\alpha$.
7 Application of System 8
System 8 can be used to compute the $(n+3)^{th}$ Dedekind number ($D(n+3)$) based on intervals in the space of antichains over $n$ elements.
Theorem 7.1
Let $n\geq 0$ be a natural number, $N=\{1,..,N\}$, $\mathcal{D}_{n}$ the lattice of antichains over $N$. We have
$D(n+3)=\sum_{\begin{subarray}{c}\alpha\in\mathcal{D}_{n}\\ \beta_{1},\beta_{2},\beta_{3}\in[\alpha,\top]\\ \gamma\in[\beta_{1}\vee\beta_{2}\vee\beta_{3},\top]\end{subarray}}P_{3}(\alpha,\beta_{1},\beta_{2},\beta_{3})\times|[\bot,\alpha]||[\beta_{1},\gamma]||\beta_{2},\gamma]||[\beta_{3},\gamma]$
  • Proof
    We use the decomposition for $\eta\in\mathcal{D}_{n+3}$ in terms of elements of $\mathcal{D}_{n}$:
    $\eta=$ $\eta_{0}\times\{123\}$
    $\vee\eta_{1}\times\{23\}\vee\eta_{2}\times\{13\}\vee\eta_{3}\times\{12\}$
    $\vee\eta_{12}\times\{3\}\vee\eta_{13}\times\{2\}\vee\eta_{23}\vee\{1\}$
    $\vee\eta_{123}\times\{0\}$
    were sets such as $\{1,3\}$ are denoted by $13$ and $\bot$ by 00 as usual. The decomposition is unique iff $A\subseteq B\Rightarrow\eta_{A}\leq\eta_{B}$. We identify $\eta_{1},\eta_{2},\eta_{3}$ with the variables $\chi_{1},\chi_{2},\chi_{3}$ in System 8. Any solution of System 8 for specific values of $\alpha,\beta_{1},\beta_{2},\beta_{3}$ will now correspond to antichains $\eta$ for which $\eta_{0}\leq\alpha$, $\eta_{123}\geq\gamma$ for any $\gamma\geq\beta_{1}\vee\beta_{2}\vee\beta_{3}$ and $\eta_{12}\in[\beta_{3},\gamma],\eta_{13}\in[\beta_{2},\gamma],\eta_{23}\in[\beta_{1},\gamma]$. $\qed$
  • Example
[…]
  • Example
[…]
$\alpha$ $\beta_{1}$ $\beta_{2}$ $\beta_{3}$Eq $P_{3}$ $\sum_{\gamma\in[\beta_{1}\vee\beta_{2}\vee\beta_{3},\{0\}]}|[\bot,\alpha]||[\beta_{1},\gamma]||[\beta_{2},\gamma]||[\beta_{3},\gamma]|$ Tot
$\bot$ $\bot$ $\bot$ $\bot$ 1 1 1 + 8 9
$\bot$ $\{0\}$$\{0\}$ $\bot$ 3 1 2 6
$\bot$ $\{0\}$$\{0\}$$\{0\}$ 1 3 1 3
$\{0\}$$\{0\}$$\{0\}$$\{0\}$ 1 1 2 2
$\sum$ 20
Table 3: Simplest example of System 8, the result is D(3)
$\alpha$ $\beta_{1}$ $\beta_{2}$ $\beta_{3}$Eq $P_{3}$ $|[\bot,\alpha]|\times\sum_{\gamma\in[\beta_{1}\vee\beta_{2}\vee\beta_{3},\{1\}]}|[\beta_{1},\gamma]||[\beta_{2},\gamma]||[\beta_{3},\gamma]|$ Tot
$\bot$ $\bot$ $\bot$ $\bot$ 1 1 1 + 8 + 27 = 36 36
$\bot$ $\{0\}$$\{0\}$ $\bot$ 3 1 2 + 12 = 14 42
$\bot$ $\{0\}$$\{0\}$$\{0\}$ 1 3 1 + 8 = 9 27
$\bot$ $\{1\}$$\{1\}$ $\bot$ 3 1 3 9
$\bot$ $\{1\}$$\{1\}$$\{0\}$ 3 2 2 12
$\bot$ $\{1\}$$\{1\}$$\{1\}$ 1 3 1 3
$\{0\}$$\{0\}$$\{0\}$$\{0\}$ 1 1 2 $\times$ (1 + 8) = 18 18
$\{0\}$$\{1\}$$\{1\}$$\{0\}$ 3 1 2$\times$2 = 4 12
$\{0\}$$\{1\}$$\{1\}$$\{1\}$ 1 3 2$\times$1 = 2 6
$\{1\}$$\{1\}$$\{1\}$$\{1\}$ 1 1 3$\times$1 = 2 3
$\sum$ 168
Table 4: One but simplest example of System 8, the result is D(4)
8 A system of eight equations
Let $\alpha_{i},\beta_{i},i\not=j\in\{1,2,3,4\}$ be eight antichains with $\alpha_{i}\leq\beta_{j}$. Consider the set of eight equations in six variables $\chi_{12},\chi_{13},\chi_{14},\chi_{23},\chi_{24},\chi_{34}$:
$\chi_{12}\wedge\chi_{13}\wedge\chi_{14}=\alpha_{1}$ (12)
$\chi_{12}\wedge\chi_{23}\wedge\chi_{24}=\alpha_{2}$ (13)
$\chi_{13}\wedge\chi_{23}\wedge\chi_{34}=\alpha_{3}$ (14)
$\chi_{14}\wedge\chi_{24}\wedge\chi_{34}=\alpha_{4}$ (15)
$\chi_{23}\vee\chi_{24}\vee\chi_{34}=\beta_{1}$ (16)
$\chi_{13}\vee\chi_{14}\vee\chi_{34}=\beta_{2}$ (17)
$\chi_{12}\vee\chi_{14}\vee\chi_{24}=\beta_{3}$ (18)
$\chi_{12}\vee\chi_{13}\vee\chi_{23}=\beta_{4}$ (19)
  • Notation
    We denote by $S_{\alpha}=\alpha_{1}\cup\alpha_{2}\cup\alpha_{3}\cup\alpha_{4},S_{\beta}=\beta_{1}\cup\beta_{2}\cup\beta_{3}\cup\beta_{4},$ the unions of the antichains $\alpha_{i}$ and $\beta_{i}$.
Lemma 8.1
System 12 possibly accepts solutions including sets not in $S_{\alpha}\cup S_{\beta}$.
  • Proof
    We give a counterexample. Let $\beta_{1}=\beta_{4}=\{1234\},\beta_{2}=\beta_{3}=\{123\}$, $\alpha_{1}=\alpha_{2}=\alpha_{3}=\alpha_{4}=\{1\}$ A set of solutions of System 12 is given by $\chi_{12}=\chi_{13}=\chi_{24}=\{1\},\chi_{23}=\{1234\},\chi_{14}=\{123\},\chi_{34}\in[\{1\},\{123\}]$. $\chi_{34}$ can contain sets $3,2$ 3 , 13 , 23 which are not in $S_{\alpha}\cup S_{\beta}$. $\qed$
Due to Lemma 8.1, the number of solutions of System 12 cannot be expected to be computable in polynomial time.
  • Notation
    We denote by $P_{a}(,\alpha_{2},\alpha_{3},\alpha_{4},\beta_{1},\beta_{2},\beta_{3},\beta_{4})$ the number of solutions of System 12 for the given antichains $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ and $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ on the righthand sides of the equations.
9 Application of System 12
System 12 can be used to compute the $(n+4)^{th}$ Dedekind number ($D(n+4)$) based on intervals in the space of antichains over $n$ elements.
Theorem 9.1
Let $n\geq 0$ be a natural number, $N=\{1,..,N\}$, $\mathcal{D}_{n}$ the lattice of antichains over $N$. We have
$D(n+4)=$ $\sum_{\begin{subarray}{c}\alpha_{1}\leq\beta_{2}\wedge\beta_{3}\wedge\beta_{4}\\ \alpha_{2}\leq\beta_{1}\wedge\beta_{3}\wedge\beta_{4}\\ \alpha_{3}\leq\beta_{1}\wedge\beta_{2}\wedge\beta_{4}\\ \alpha_{4}\leq\beta_{1}\wedge\beta_{2}\wedge\beta_{3}\\ \gamma\leq\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3}\wedge\alpha_{4}\\ \delta\geq\beta_{1}\vee\beta_{2}\vee\beta_{3}\vee\beta_{4}\end{subarray}}$
$P_{4}(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\beta_{1},\beta_{2},\beta_{3},\beta_{4})$
$\times$ $|[\gamma,\alpha_{1}]||[\gamma,\alpha_{2}]||[\gamma,\alpha_{3}]||[\gamma,\alpha_{4}]|$
$\times$ $|[\beta_{1},\delta]||[\beta_{2},\delta]||[\beta_{3},\delta]||[\beta_{4},\delta]|$
  • Proof
    We use the decomposition for $\eta\in\mathcal{D}_{n+4}$ in terms of elements of $\mathcal{D}_{n}$:
    $\eta=$ $\eta_{0}\times\{1234\}$
    $\vee\eta_{1}\times\{234\}\vee\eta_{2}\times\{134\}\vee\eta_{3}\times\{124\}\vee\eta_{4}\times{123}$
    $\vee\eta_{12}\times\{34\}\vee\eta_{13}\times\{24\}\vee\eta_{14}\times\{23\}\vee\eta_{23}\times{14}\vee\eta_{24}\times\{13\}\vee\eta_{34}\times{12}$
    $\vee\eta_{123}\times\{4\}\vee\eta_{124}\times\{3\}\vee\eta_{134}\times\{2\}\vee\eta_{234}\times\{1\}$
    $\vee\eta_{1234}\times\{0\}$
    The decomposition is unique iff $A\subseteq B\Rightarrow\eta_{A}\leq\eta_{B}$. We identify $\eta_{12},\eta_{13},\eta_{14},\eta_{23},\eta_{24},\eta_{34}$ with the variables $\chi_{12},\chi_{13},\chi_{14},\chi_{23},\chi_{24},\chi_{34}$ in System 12. Any solution of System 12 for specific values of $\alpha_{i},\beta_{i}$ will now correspond to antichains $\eta$ for which $\eta_{0}=\gamma\leq\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3}\wedge\alpha_{4}$, $\eta_{1234}=\delta\geq\beta_{1}\vee\beta_{2}\vee\beta_{3}\vee\beta_{4}$, $\eta_{i}\in[\gamma,\alpha_{i}],\eta_{jkl}\in[\beta_{i},\delta]$ for $\{i,j,k,l\}=\{1,2,3,4\}$. $\qed$
10 A system of seven equations
Let $\alpha\leq\gamma_{12},\gamma_{13},\gamma_{14},\gamma_{23},\gamma_{24},\gamma_{34}\leq\beta$ be seven antichains. Consider the set of seven equations in four variables $\chi_{1},\chi_{2},\chi_{3},\chi_{4}$:
$\chi_{1}\wedge\chi_{2}\wedge\chi_{3}\wedge\chi_{4}=\alpha$ (20)
$\chi_{1}\vee\chi_{2}=\gamma_{12}$ (21)
$\chi_{1}\vee\chi_{3}=\gamma_{13}$ (22)
$\chi_{1}\vee\chi_{4}=\gamma_{14}$ (23)
$\chi_{2}\vee\chi_{3}=\gamma_{23}$ (24)
$\chi_{2}\vee\chi_{4}=\gamma_{24}$ (25)
$\chi_{3}\vee\chi_{4}=\gamma_{34}$ (26)
We will use the notation $\alpha=\alpha_{1}\cup\alpha_{2}\cup\alpha_{3}\cup\alpha_{4}$.
Lemma 10.1
In any solution of System 20, for all $i\in\{1,2,3,4\}$, we have $\chi_{i}=\alpha\vee\chi^{\prime}_{i}$ with $\chi^{\prime}_{i}\subseteq\gamma_{ij}-\alpha$ for all $j\not=i$.
  • Proof
    Suppose $X\in\chi^{\prime}_{1},X\not\in\gamma_{1j}-\alpha$ for all $j\not=i$. Due to equations 21,22 and 23, each of $\chi_{2},\chi_{3},\chi_{4}$ must contain a set $X_{j}\in\gamma_{1j}$ such that $X\subsetneq X_{j}$. As a consequence, $\alpha=\chi_{1}\wedge\chi_{2}\wedge\chi_{3}\wedge\chi_{4}\geq\{X\}$ and hence $X\in\alpha$, a contradiction. $\qed$
Lemma 10.2
In any solution of System 20, for all $i\in\{1,2,3,4\}$, $\chi_{i}=\alpha\vee\chi^{\prime}_{i}$ with $\chi^{\prime}_{i}\subseteq\cap_{j\not=i}\gamma_{ij}-\alpha$.
  • Proof
    This follows from Lemma 10.1 and the equations 21, 22 and 23.
Lemma 10.3
In any solution of System 20, for all $i\in\{1,2,3,4\}$, any set in $\cap_{j\not=i}\gamma_{ij}-\alpha$ is contained in $\chi_{i}$ or $\cap_{j\not=i}\chi_{j}$ and not in both.
  • Proof
    A set $X\in\cap_{j\not=i}\gamma_{ij}-\alpha$ cannot be in all $\chi_{i}$ due to equation 20. The two possibilities are the only ways to satisfy equations 21 - 26. $\qed$
Lemma 10.4
In any solution of System 20, for all $i\not=j\in\{1,2,3,4\}$, if two sets $X,Y\in\cap_{k\not=i}\gamma_{ik}-\alpha$ satisfy $\{X\cap Y\}\not\leq\alpha$, both sets must be together, either in $\chi_{i}$ or in $\cap_{j\not=i}\chi_{j}$.
  • Proof
    If $X$ were in $\chi_{i}$ and $Y$ in $\cap_{j\not=i}\chi_{j}$, equation 20 would be violated. $\qed$
  • Definition
    Let $i\in\{1,2,3,4\}$. Two sets $X,Y\in\cap_{k\not=i}\gamma_{ik}-\alpha$ are i-connected if $\{X\cap Y\}\not\leq\alpha$ or there is a sequence $X,Z_{1},\ldots,Z_{k},Y$ of sets in $\cap_{j\not=i}\chi_{j}$ in which any two consecutive sets are i-connected. The resulting graph is called the i-connected graph (i-g).
  • Notation
    For a graph g, $C_{g}$ denotes the number of connected components of the graph.
Lemma 10.5
In any solution of System 20, for all $i\in\{1,2,3,4\}$, the maximum number of possible assignments of sets from $\cap_{k\not=i}\gamma_{ik}-\alpha$ is $2^{C_{i-g}}$.
  • Proof
    Follows from Lemma 10.4 and definition 10$\qed$
Lemma 10.6
In any solution of System 20, for all $i\not=j\in\{1,2,3,4\}$, if $g_{i}$ is an i-connected component and $g_{j}$ is a j-connected component such that $g_{i}$ and $g_{j}$ are connected as well, at most one of $g_{i}$ or $g_{j}$ cannot be assigned to $\chi_{i}$ or $\chi_{j}$ and the number of possible combined assignments is 3.
  • Proof
    If both were assigned three times, the connected sets in $g_{i}$ and $g_{j}$ would together cover the four $\chi_{k}$ and violate equation 20$\qed$
  • Definition
    We denote the number of solutions to System 20 as $P_{4}(\alpha,\gamma_{12},\gamma_{13},\gamma_{14},\gamma_{23},\gamma_{24},\gamma_{34})$.
The Lemma’s above allow to compute $P_{4}(\alpha,\gamma_{12},\gamma_{13},\gamma_{14},\gamma_{23},\gamma_{24},\gamma_{34})$ for given $\alpha$ and $\gamma_{ij}$.
11 Application of System 20
System 20 can be used to compute the $(n+4)^{th}$ Dedekind number ($D(n+4)$) based on intervals in the space of antichains over $n$ elements.
Theorem 11.1
Let $n\geq 0$ be a natural number, $N=\{1,..,N\}$, $\mathcal{D}_{n}$ the lattice of antichains over $N$. We have
$D(n+4)=$ $\sum_{\begin{subarray}{c}\alpha\leq\gamma_{12}\wedge\gamma_{13}\wedge\gamma_{14}\wedge\gamma_{23}\wedge\gamma_{24}\wedge\gamma_{34}\\ \beta\geq\delta_{12}\vee\delta_{13}\vee\delta_{14}\vee\delta_{23}\vee\delta_{24}\vee\delta_{34}\\ \gamma_{ij}\leq\delta_{ij}\end{subarray}}$
$P_{4}(\alpha,\gamma_{12},\gamma_{13},\gamma_{14},\gamma_{23},\gamma_{24},\gamma_{34})$
$\times$ $P_{4}(\tilde{\beta},\widetilde{\delta_{12}},\widetilde{\delta_{13}},\widetilde{\delta_{14}},\widetilde{\delta_{23}},\widetilde{\delta_{24}},\widetilde{\delta_{34}})$
[…]
$\times$ $|[\gamma_{12},\delta_{12}]|\times|[\gamma_{13},\delta_{13}]|\times|[\gamma_{14},\delta_{14}]|\times|[\gamma_{23},\delta_{23}]|\times|[\gamma_{24},\delta_{24}]|\times|[\gamma_{34},\delta_{34}]|$
$\times$ $|[\beta,\{N\}]|$
  • Proof
    We use the decomposition for $\eta\in\mathcal{D}_{n+4}$ in terms of elements of $\mathcal{D}_{n}$:
    $\eta=$ $\eta_{0}\times\{1234\}$
    $\vee\eta_{1}\times\{234\}\vee\eta_{2}\times\{134\}\vee\eta_{3}\times\{124\}\vee\eta_{4}\times{123}$
    $\vee\eta_{12}\times\{34\}\vee\eta_{13}\times\{24\}\vee\eta_{14}\times\{23\}\vee\eta_{23}\times{14}\vee\eta_{24}\times\{13\}\vee\eta_{34}\times{12}$
    $\vee\eta_{123}\times\{4\}\vee\eta_{124}\times\{3\}\vee\eta_{134}\times\{2\}\vee\eta_{234}\times\{1\}$
    $\vee\eta_{1234}\times\{0\}$
    The decomposition is unique iff $A\subseteq B\Rightarrow\eta_{A}\leq\eta_{B}$. We identify $\eta_{1},\eta_{2},\eta_{3},\eta_{4}$ with the variables $\chi_{1},\chi_{2},\chi_{3},\chi_{4}$ in System 20, and $\eta_{123},\eta_{124},\eta_{134},\eta_{234}$ with the duals of the variables $\upsilon_{1},\upsilon_{2},\upsilon_{3},\upsilon_{4}$ Any combined solution of the two realisations of System 20 for specific values of $\alpha,\beta$ and $\gamma_{ij}$ will now correspond to antichains $\eta$ for which $\eta_{0}\leq\alpha$, $\eta_{1234}\geq\beta$, $\gamma_{ij}\leq\eta_{ij}\leq\delta_{ij}$. The product of the two $P_{4}$ coefficients counts the number of combined solutions for the $\chi_{i},\upsilon_{i}$$\qed$
12 Conclusion
We presented a number of systems of equations in the free distributive lattice of antichains and introduced a technique to count the number of solutions. The systems of equations were chosen to be applicable to the problem of counting the number of antichains of subsets of a finite set of $n$ elements, the $n^{t}h$ Dedekind number $D(n)$ which is the size of the antichain set $\mathcal{D}(n)$. We could apply our technique to reduce $D(n+2)$ to a sum over products of interval sizes in $\mathcal{D}(n)$, which was used in our computation of the ninth Dedekind number. We showed that the technique can be applied for similar reductions of $D(n+3)$ and $D(n+4)$.