### v3

# Solving systems of equations on antichains for the computation of the ninth Dedekind Number

Abstract
We study four systems of equations, together with a way to count the number of solutions. One of the results was used in the recent computation of D(9), the others have potential to speed up existing techniques in the future.
keywords:
Dedekind numbers, antichains, algebra of distributive lattices, posets, intervals
journal: Journal of  Templates
1 Introduction
Calculating Dedekind Numbers (OEIS series A000372 [1]) has always followed progress in both the Mathemathical domain, and the Computational Implementation Domain. Computing Subsequent Numbers is known as ”Dedekind’s Problem”. Since the year 1991 and up until March 2023, the highest known term in this sequence was D(8), computed by Doug Wiedemann [2] on a Cray 2 supercomputer. In April 2023 two independent computations of D(9) were announced (see [3, 4]). The computations happened completely independently, were based on different methodologies and relied on very different hardware. Christian Jäkel ([3]) used a GPU infrastructure and built on a matrix formulation, which allowed to use the advanced techniques and highly developed software for numerical analysis. In our computation (Van Hirtum et al. [4]), we, together with our co-authors, relied on the advanced super computing infrastructure at the PC2 centre at Paderborn University, using cutting edge FPGA hardware and an ingenious implementation of the so-called P-coefficient formula [5]. At the end of the latter computation, some uncertainty due to the semi-experimental setting of the FPGA infrastructure was identified. Although we obtained the result on March 8, it was only when the result of Jäkel came out, confirming our result, that we could be certain. An extensive report on this computational process is in [4], concentrating on the very important computational aspects and the ingenious implementation which brought new insights for FPGA technology. The present paper elaborates on the mathematical aspects, with a discussion on the P-coefficient and possible extensions.
2 Definitions and notations
2.1 The Dedekind number
For any finite positive number $n\geq 0$, the n-th Dedekind number counts the number of monotone Boolean functions, or equivalently, the number of antichains, on the subsets of a finite set $\mathcal{A}=\{1,\dots,n\}$. To derive the properties in this paper, we will use the concept of an antichain over $\mathcal{A}$, i.e. sets of pairwise non-inclusive subsets of $\mathcal{A}$. The set of all antichains over $\mathcal{A}$ will be denoted by $\mathcal{D}_{n}$, its size, the $n^{th}$ Dedekind number, will be denoted by $D(n)$.
The set of permutations of the elements of base set $\mathcal{A}$ generates an equivalence relation on $\mathcal{D}_{n}$. The set of equivalence classes of this relation is denoted by $\mathcal{R}_{n}$ and the number of such equivalence classes is denoted by $R(n)$.
The known Dedekind numbers are shown in Table 1, including the result of our computation. Table 2 shows the known numbers $R(n)$ of equivalence classes of monotone Boolean functions under permutation of the elements of the base set. Note that the last result dates from 2023.
D(0) 2 Dedekind (1897)
D(1) 3 Dedekind (1897)
D(2) 6 Dedekind (1897)
D(3) 20 Dedekind (1897)
D(4) 168 Dedekind (1897)
D(5) 7581 Church (1940)
D(6) 7828354 Ward (1946)
D(7) 2414682040998 Church (1965)
D(8) 56130437228687557907788 Wiedemann (1991)
D(9) 286386577668298411128469151667598498812366 Jäkel, Van Hirtum et al. (2023)
Table 1: Known Dedekind Numbers, OEIS series A000372 [1], including our result.
R(0) 2
R(1) 3
R(2) 5
R(3) 10
R(4) 30
R(5) 210
R(6) 16353
R(7) 490013148 Tamon Stephen & Timothy Yusun (2014) [6]
R(8) 1392195548889993358 Bartelomiej Pawelski (2021) [7]
R(9) 789204635842035040527740846300252680 Bartelomiej Pawelski (2023)
Table 2: Known Equivalence Class Counts, OEIS series A003182 [8]
2.2 Notation
Let $n\in\mathbb{N},\mathcal{A}=\{1,\dots,n\},2^{\mathcal{A}}=\{X\subseteq\mathcal{A}\},\mathbb{B}=\{false,true\}$.
  • Definition
    A Boolean function $f:2^{\mathcal{A}}\rightarrow\mathbb{B}$ is monotone iff
    $\forall X,Y\in 2^{\mathcal{A}}:X\subseteq Y\Rightarrow f(Y)\Rightarrow f(X)$
Note that a monotone Boolean function is completely defined by the set of sets which are maximal among the sets for which the function value is true.
  • Definition
    A set $X\in 2^{\mathcal{A}}$ is a maximal set of a monotone Boolean function $f$ iff
    $f(X)=true\ and\ \forall Y\in 2^{\mathcal{A}}:X\subsetneq Y\Rightarrow f(Y)=false$
For any monotone Boolean function, no two of its maximal sets include each other. A set of sets with this property is called an antichain:
  • Definition
    A set of sets $\sigma\subseteq 2^{\mathcal{A}}$ is an antichain iff
    $\forall X,Y\in\sigma:X\not\subseteq Y\ and\ Y\not\subseteq X$
In what follows, we will represent antichains by letters from the Greek alphabet $\alpha,\beta,\ldots$. The elements of the set $\mathcal{A}$ will be denoted by Latin minuscules $a,b,c,\ldots$ or by digits $1\ldots 9$. For subsets, unless there is a possibility for confusion, we will use the abbreviation $abc\equiv\{a,b,c\},123\equiv\{1,2,3\},0\equiv\emptyset$ or refer to them by by Latin majuscules $X,Y,\ldots$.
  • Definition
    An antichain $\alpha$ is less or equal than an antichain $\beta$ iff each set in $\alpha$ is included in at least one set of $\beta$:
    $\alpha\leq\beta\Leftrightarrow\forall X\in\alpha:\exists Y\in\beta:X\subseteq Y$
  • Notation
    We denote by $\bot$ and $\top$ the smallest, respectively the largest, element of $\mathcal{D}_{n}$:
    {},$\bot\equiv\{\},\top\equiv\{12\ldots n\}$ (1)
    $\forall\alpha\in\mathcal{D}_{n}:\bot\leq\alpha\leq\top$ (2)
  • Definition
    For any two elements $\alpha,\beta\in\mathcal{D}_{n}$, the interval with bottom $\alpha$ and top $\beta$ is the set of antichains $\chi$ satisfying $\alpha\leq\chi\leq\beta$:
    $\forall\alpha,\beta\in\mathcal{D}_{n}:[\alpha,\beta]=\{\chi\in\mathcal{D}_{n}:\alpha\leq\chi\leq\beta\}$
  • Definition
    For $\alpha,\beta\in\mathcal{D}_{n}$, the join $\alpha\vee\beta$ and the meet $\alpha\wedge\beta$ are the antichains defined by
    $\alpha\vee\beta=max(\{X\in\alpha\cup\beta\})$ (3)
    $\alpha\wedge\beta=max(\{X\cap Y|X\in\alpha,Y\in\beta\})$ (4)
Where $max$ denotes the maximum w.r.t. inclusion of a set of subsets of $\mathcal{A}$, i.e.
$S\subset 2^{\mathcal{A}}\Rightarrow max(S)=\{X\in S:\forall Y\in S:X\not\subsetneq Y\}$
As an antichain is a set of sets, we use the operations intersection ()($\cap$)( ∩ ), union ()($\cup$)( ∪ ), subset ()($\subseteq$)( ⊆ ) and difference ()(-)( - ) to work on antichains. These should not be confused with the join ()($\vee$)( ∨ ), meet ()($\wedge$)( ∧ ), and less or equal ()($\leq$)( ≤ ) operations in the partial order we have just introduced. We will sometimes use the span of an antichain which is the union of its elements:
  • Definition
    $\forall\alpha\in\mathcal{D}_{n}:sp(\alpha)=\cup_{X\in\alpha}X$
A last operator on two antichains is the direct product $x$ defined for any two antichains $\alpha$ and $\beta$ for which the span does not overlap:
  • Definition
    $\forall\alpha,\beta\in\mathcal{D}_{n},sp(\alpha)\cap sp(\beta)=\emptyset:\alpha\times\beta=\{X\cup Y|X\in\alpha,Y\in\beta\}$
Finally, in the formulas below, a number defined for each pair $\alpha\leq\beta\in\mathcal{D}_{n}$ plays an important role. We refer to this number as the connector number $C_{\alpha,\beta}$. It counts the number of connected components of a graph with sets in $\beta$ as vertices and edges between two sets if the intersection is not dominated by $\alpha$ (exact definition follows).
3 Related Work
The work in this article builds upon three papers by Patrick De Causmaecker et al. [9, 5, 10]. Recently, a number of interesting results were published by Bartolomiej Pawelski, e.g. the recent computation of $R(9)$ [11].
Finally of course we want to mention the work of Christian Jäkel, who computed the 9th Dedekind Number at about the same time as us, but using different formulas and different hardware. [3].
Central to our computation was a result on the number of solutions of the set of two equations
$\chi\vee\upsilon=\beta$ (5)
$\chi\wedge\upsilon=\alpha$ (6)
for $\chi,\upsilon\in D_{n}$ is given by $2^{C_{\alpha,\beta}}$. This is called the P-Coefficient also denoted as PCoeff [5, 10]. Below we will elaborate on this result and study its extension to larger systems of equations.
In the following sections, we derive various P-coefficient formulas.
4 A system of two equations
Let $\alpha\leq\beta$ be two antichains. Consider the set of equations, hereafter refered to as System 4
$\chi\wedge\upsilon=\alpha$ (7)
$\chi\vee\upsilon=\beta$ (8)
It is clear that the system has no solution if $\alpha\not\leq\beta$. The following Lemma states that in any solution $\chi,\upsilon$, only the sets in the righthand sides of System 4 can occur. It significantly simplifies the construction as well as the counting of solutions of System 4 .
Lemma 4.1
In any solution of System 4, $\chi=\alpha\vee\chi_{1},\upsilon=\alpha\vee\upsilon_{1}$ with $\chi_{1}\subseteq\beta-\alpha,\upsilon_{1}\subseteq\beta-\alpha$.
  • Proof
    Suppose $X\in\chi_{1},X\not\in\beta$. Due to equation 8, there must be a set $X^{\prime}\in\upsilon$ such that $X\subseteq X^{\prime}$ and $X^{\prime}\in\beta$. Due to equation 7, this implies $X\in\alpha$ in contradiction with $\chi_{1}\subseteq\beta-\alpha$. Similar arguments work for $\upsilon_{1}$. $\qed$
The following lemma states that any set satisfying the condition of Lemma 4.1 occurs in exactly one of the two antichain variables $\chi$ or $\upsilon$.
Lemma 4.2
For any set $A\in\beta-\alpha$, we have that $A\in\chi-\upsilon$ or $A\in\upsilon-\chi$.
  • Proof
    From equation 8, it follows that $A$ must be in at least one of $\chi$ or $\upsilon$. From the same equation it follows that no set dominating $A$ can be in any of the two. If $A$ were in both $\chi$ and $\upsilon$, $A\in\chi\wedge\upsilon$ and equation 7 would be violated. $\qed$
Lemma 4.3
Let $A,B\in\beta-\alpha$, such that $\{A\cap B\}\not\leq\alpha$. We have
$A\in\chi\Leftrightarrow B\in\chi$
$A\in\upsilon\Leftrightarrow B\in\upsilon$
  • Proof
    Due to Lemma 4.2, $A$ and $B$ are each in exactly one of $\chi$ and $\upsilon$. If they were each in a different one, we would have $\{A\cap B\}\leq\chi\wedge\upsilon=\alpha$ due to equation 7. A contradiction. $\qed$
  • Definition
    Given two antichains $\alpha\leq\beta$, two sets $A,B\in\beta-\alpha$ are said to be directly connected with respect to $\alpha$ iff $\{A\cap B\}\not\leq\alpha$.
  • Definition
    Two sets $A,B\in\beta-\alpha$ are said to be connected w.r.t. $\alpha$ iff $A$ and $B$ are directly connected w.r.t. $\alpha$ or there exist $Z_{1},Z_{2},\ldots,Z_{x}\in\beta$, $x\geq 1$, such that $A$ is directly connected with $Z_{1}$, $B$ is directly connected with $Z_{x}$ and each $Z_{i}$ is directly connected to $Z_{i+1}$ for $i\in\{1,\ldots,x-1\}$, all w.r.t. $\alpha$.
  • Definition
    For two antichains $\alpha\leq\beta$, the connection graph $C_{\alpha}(\beta)$ is the graph with the sets of $\beta$ as vertices and edges
    $E=\{(A,B)|A,B\in\beta,A,B\text{ are connected w.r.t. }\alpha\}$
Theorem 4.4
Let $\alpha\leq\beta$ be two antichains, $(\chi_{0},\upsilon_{0})$ a partition of $\beta-\alpha$ such that any two sets $A,B\in\beta-\alpha$ that are connected w.r.t. $\alpha$ are in the same component of the partition. With the partition $(\chi_{0},\upsilon_{0})$ corresponds exactly one solution of System 4 and any solution of System 4 corresponds with exactly one such partition.
  • Proof
    Due to Lemma 4.1, any solution is determined by two subsets of $\beta-\alpha$. Due to lemma 4.3, any two directly connected sets $A,B\in\beta$ must come in the same antichain $\chi$ or $\upsilon$. By transitive closure, this must be true for any two connected sets $A,B\in\beta$ and hence for all sets in any connected component of $C_{\alpha}(\beta)$. Given a partition $(\chi_{0},\upsilon_{0})$ of $\beta-\alpha$ satisfying the criteria of the theorem, we arrive at a solution $\chi=\chi_{0}\vee\alpha,\upsilon=\upsilon_{0}\vee\alpha$. The reverse follows by identifying $\chi_{0}=\chi-\alpha,\upsilon_{0}=\upsilon-\alpha$. $\qed$
Corollary 4.5
The number of solutions of System 4 is given by
$2^{C_{\alpha,\beta}}$
where $C_{\alpha,\beta}$ is the number of connected components of the connection graph $C_{\alpha}(\beta)$.
  • Proof
    The number of partitions in two sets of the connected components of $C_{\alpha}(\beta)$ is given by $2^{C_{\alpha,\beta}}$. As was proven in Theorem 4.4, this is the number of solutions of System 4.
Corollary 4.6
The number of solutions to system 4 can be computed in polynomial time on the number of sets in $\beta-\alpha$.
  • Proof
    By Theorem 4.4, the problem is reduced to counting the connected components in an undirected graph with sets in $\beta-\alpha$ as vertices. This can be done in $O(V+E)$ time where $V$ is the number of vertices and $E$ is the number of edges. $\qed$
¡ltx:note¿T¡/ltx:note¿he number $2^{C_{\alpha,\beta}}$ is referred to as the P-coefficient in [4].
5 Application of System 4
System 4 can be used to compute the $(n+2)^{th}$ Dedekind number ($D(n+2)$) based on intervals in the space of antichains over $n$ elements.
Theorem 5.1
Let $n\geq 0$ be a natural number, $N=\{1,..,N\}$, $\mathcal{A}_{n}$ the lattice of antichains over $N$. We have
$D(n+2)=\sum_{\alpha\leq\beta\in\mathcal{D}_{n}}2^{C_{\alpha,\beta}}|[\bot,\alpha]||[\beta,\top]|$
  • Proof
    We use the decomposition for $\eta\in\mathcal{D}_{n+2}$ in terms of elements $\eta_{X}$ of $\mathcal{D}_{n}$:
    $\eta=$ $\eta_{0}\times\{12\}$
    $\vee\eta_{1}\times\{1\}\vee\eta_{2}\times\{2\}$
    $\vee\eta_{12}\times\{0\}$
    were sets such as $\{1,2\}$ are denoted by $12$ and $\bot$ by 00 as usual. The decomposition is unique iff $A\subseteq B\Rightarrow\eta_{A}\leq\eta_{B}$. We identify $\eta_{1},\eta_{2}$ with the variables $\chi,\upsilon$ in System 4. Any solution of System 4 for specific values of $\alpha,\beta$ will now correspond to exactly one antichain $\eta$ for which $\eta_{0}\in[\bot,\alpha]$, $\eta_{12}\in[\beta,\top]$. $\qed$
Corollary 5.2
$D(n+2)$ can be computed as a sum of $D(n-1)$ terms, each of which is the product of the sizes of two intervals in $\mathcal{D}_{n}$.
  • Proof
    The condition $\alpha\leq\beta$ for $\alpha,\beta\in\mathcal{D}_{n}$ allows exactly $D(n-1)$ terms.$\qed$
This improves on the number of terms used in the computation of $D(8)$ by Wiedemann [2], because without the condition $\alpha\leq\beta$, the number of terms is $D(n-1)^{2}$. Theorem 5.1 and its corollary 5.2 were the basis for our computation of $D(9)$ [4]. Of course, as Wiedemann did as well, we had to reduce the number of terms further using permutation symmetry.
6 General case
Let $\alpha,\{\beta_{ij}|1\leq i\leq r,1\leq j\leq r,i\not=j\}\}$ be a set of antichains for some integer $r>1$. We will assume $\beta_{ij}=\beta_{ji}$ and will not distinguish between them. We consider the following set of of $1+\frac{r(r+1)}{2}$ equations in $r$ variables $\chi_{k},1\leq k\leq r$, hereafter referred to as System 6:
$\bigwedge_{i=1}^{r}\chi_{i}$ $=\alpha$ (9)
$\chi_{k}\vee\chi_{l}$ =$\beta_{kl}\ \ \ \ \ (1\leq k\leq r,1\leq l\leq r,k\not=l)$ (10)
  • Notation
    We will use $\beta=\cup_{k\not=l}\beta_{kl}$, $\gamma=\beta-\alpha$.
Note that for System 6 to have a solution, no set in $\gamma$ can be dominated by $\alpha$. Given equation 10, for any solution of System 6, for $i,j\in\{1,\ldots,r\}$, each set in $\beta_{ij}$ must be in $\chi_{i}$ or $\chi_{j}$.
The following Lemma proves the reverse is also true: the only sets occurring in a solution of System 6 are sets in $\beta$.
Lemma 6.1
In any solution of System 6, $\chi_{k}\subseteq\beta$ for $k\in\{1,\ldots,r\}$.
  • Proof
    Let $X\in\chi_{k},X\not\in\gamma$. According to the equation 10 $\chi_{k}\vee\chi_{i}=\beta_{ki}$ for any $i\not=k$, there must be a set $Y_{i}\in\beta_{ki}$ s.t. $X\subsetneq Y_{i}$. To satisfy the equation, we must have $Y_{i}\in\chi_{i}$. According to equation 9, we find $\alpha=\chi_{k}\wedge(\wedge_{i\not=k}\chi_{i})\geq\{X\}$ and hence $X\in\alpha$. $\qed$
  • Definition
    For any antichain $\delta$, a set $X$ is said to be dominated by $\delta$ if $\{X\}\leq\delta$.
  • Definition
    A set $X$ is said to be dominating in a set of sets $\rho$ if $X\in\rho$ and there is no set $Y\in\rho$ s.t. $X\subsetneq Y$.
We find that for any $X$ not dominated by at least one of $\beta_{ij}$, the set of indices $s$ s.t. $X\in\chi_{s}$ is fixed.
Lemma 6.2
For any set $X$ dominating in $\gamma$, if there is at least one $\beta_{kl}$ such that $X\not\in\beta_{kl}$, there is a set $S_{X}\subseteq\{1,\ldots,r\},S_{X}\not=\emptyset$ such that in any solution of System 6
$s\in S_{X}\Leftrightarrow\forall i\in\{1,\ldots,r\}-\{s\}:X\in\beta_{si}$
  • Proof
    $X\not\in\beta_{kl}\Rightarrow X\not\in\chi_{k}\text{ and }X\not\in\chi_{l}$. $X$ is in at least one of $\beta_{uv}$ and hence in $\chi_{u}$ or $\chi_{v}$. If $X\in\chi_{u}$ then $u\not=k,u\not=l$ and $X\in\beta_{ui}$ for any $i\in\{1,\ldots,r\}$. In particular $X\in\beta_{uk}$ and $X\in\beta_{ul}$ implying that $X$ must be in $\chi_{u}$ in any solution. The same reasoning follows if $X\in\chi_{v}$ and the lemma follows. $\qed$
Lemma 6.3
For any set $X$ dominating in $\gamma$, if $X\in\beta_{ij}$ for all $1\leq i\leq r,1\leq j\leq r,i\not=j$, in any solution of System
refsec:meetrjoin, there is exactly one $\chi_{k}$, $1\leq k\leq r$, such that $X\not\in\chi_{k}$.
  • Proof
    Suppose $X\not\in\chi_{k},X\not\in\chi_{l},k\not=l$. Then $X\not\in\beta_{kl}$, in contradiction with the condition in the lemma. Suppose $X\in\chi_{k}$ for all $1\leq k\leq r$. Then equation 9 is violated. $\qed$
  • Definition
    Two sets $X,Y\in\gamma$ are said to be directly connected w.r.t $\alpha$ if $\{X\cap Y\}\not\leq\alpha$
  • Definition
    Two sets $X,Y\in\gamma$ are said to be connected w.r.t $\alpha$ if they are directly connected w.r.t $\alpha$ or if there is a sequence $Z_{1},Z_{2},\ldots,Z_{k}$ of sets in $\gamma$ such that $X=Z_{1},Y=Z_{k}$ and $Z_{i},Z_{i+1}$ are directly connected for $1\leq i<k$.
  • Definition
    We denote by $g_{\alpha,\gamma}$ the graph with sets in $\gamma$ as vertices where there is an edge between two sets $X,Y$ iff $X$ and $Y$ are connected with respect to $\alpha$.
  • Definition
    We denote by $C_{\alpha,\gamma}$ set of connected components of $g_{\alpha,\gamma}$
As is well known, the connected components of $g_{\alpha,\gamma}$ can be found in polynomial time Ø$\O(V+E)$ where $V$ is the number of vertices in $g_{\alpha,\gamma}$, i.e. the number of sets in $\gamma-\alpha$ and $E$ the number of edges ([12]).
  • Definition
    For a connected component $c\in C_{\alpha,\gamma}$, using Lemma 6.2, its weight $w(c)$ is defined as follows:
      In case $c$ does not contain a set occurring in all $\beta_{ij}$:
      $\text{if }\cup_{X\in c}S_{X}=\{1,ldots,r\}\text{ then }w(c)=0\text{ else }w(c)=1$
      In case $c$ contains at least one set dominated by all $\beta_{ij}$:
      $w(c)=r-|\cup_{X\in c}S_{X}|$
We now prove the following
Theorem 6.4
The number of solutions of System 6 is
$P_{r}(\alpha,\beta_{ij}\text{ for }1\leq i<j\leq r)=\Pi_{c\in C_{\alpha,\gamma}}w(c)$ (11)
  • Proof
    Note that, in any solution of System 6 for each connected component $c\in\mathcal{C}_{\alpha,\gamma}$, there must be at least one $\chi_{i}$ not containing any of the sets in $c$ because if this were not the case, equation 9 would be violated. We now reduce the system set by set as follows
    1. 1.
      Copy $\gamma$ to $\gamma^{\prime}$, all $\beta_{ij}$ to $\beta^{\prime}_{ij}$. Let all $\chi_{i}=\{\}$ be empty antichains.
    2. 2.
      Choose a dominating set $X$ in $\gamma^{\prime}$ which does not occur in all $\beta^{\prime}_{ij}$. If no such set exists goto step 7
    3. 3.
      If the set $X$ does not occur in all $\beta^{\prime}_{ij}$, apply Lemma 6.2 to assign the set to the right $\chi_{i}$’s.
    4. 4.
      Remove $X$ from $\gamma^{\prime}$ and from each $\beta^{\prime}_{ij}$.
    5. 5.
      For any strict subset $Y\subsetneq X$ present in $\gamma^{\prime}$, replace each $\beta^{\prime}_{ij}$ by $\beta^{\prime}_{ij}\vee\{Y\}$.
    6. 6.
      Go to step 2
    7. 7.
      Compute the weight of each $c\in C_{\alpha,\gamma}$ using the constructed $\chi_{i}$ as partial solutions.
    The formula in the theorem now counts the number of solutions. $\qed$
7 A system of four equations
Let $\alpha\leq\beta_{23},\beta_{13},\beta_{12}$ be four antichains with $\alpha\leq\beta_{i}$ for $i\in{1,2,3}$. Consider the set of equations, herafter referred to as System 7:
$\chi_{1}\wedge\chi_{2}\wedge\chi_{3}=\alpha$ (12)
$\chi_{1}\vee\chi_{2}=\beta_{12}$ (13)
$\chi_{1}\vee\chi_{3}=\beta_{13}$ (14)
$\chi_{2}\vee\chi_{3}=\beta_{23}$ (15)
This is equivalent to the general System 6 for $r=3$. The number of solutions is given by
$P_{3}(\alpha,\beta_{12},\beta_{13},\beta_{23})=\prod_{\begin{subarray}{c}g\text{ connected component}\\ \text{ of }C_{\alpha}(\gamma)\end{subarray}}\ w(g)$
Corollary 7.1
The number of solutions of System 7 can be computed in polynomial time in the number of sets in $\beta-\alpha$.
8 Application of System 7
System 7 can be used to compute the $(n+3)^{th}$ Dedekind number ($D(n+3)$) based on intervals in the space of antichains over $n$ elements.
Theorem 8.1
Let $n\geq 0$ be a natural number, $\mathcal{A}=\{1,..,n\}$, $\mathcal{D}_{n}$ the lattice of antichains over $\mathcal{A}$. We have
$D(n+3)=\sum_{\begin{subarray}{c}\alpha\in\mathcal{D}_{n}\\ \beta_{12},\beta_{13},\beta_{23}\in[\alpha,\top]\\ \gamma\in[\beta_{12}\vee\beta_{13}\vee\beta_{23},\top]\end{subarray}}P_{3}(\alpha,\beta_{23},\beta_{13},\beta_{12})|[\bot,\alpha]||[\beta_{12},\gamma]||\beta_{13},\gamma]||[\beta_{23},\gamma]$
  • Proof
    We use the decomposition for $\eta\in\mathcal{D}_{n+3}$ in terms of elements of $\mathcal{D}_{n}$:
    $\eta=$ $\eta_{0}\times\{123\}$
    $\vee\eta_{1}\times\{23\}\vee\eta_{2}\times\{13\}\vee\eta_{3}\times\{12\}$
    $\vee\eta_{12}\times\{3\}\vee\eta_{13}\times\{2\}\vee\eta_{23}\vee\{1\}$
    $\vee\eta_{123}\times\{0\}$
    were sets such as $\{1,3\}$ are denoted by $13$ and $\bot$ by 00 as usual. The decomposition is unique iff $A\subseteq B\Rightarrow\eta_{A}\leq\eta_{B}$. We identify $\eta_{1},\eta_{2},\eta_{3}$ with the variables $\chi_{1},\chi_{2},\chi_{3}$ in System 7. Any solution of System 7 for specific values of $\alpha,\beta_{12},\beta_{13},\beta_{23}$ will now correspond to antichains $\eta$ for which $\eta_{0}\leq\alpha$, $\eta_{123}\geq\gamma$ for any $\gamma\geq\beta_{23}\vee\beta_{13}\vee\beta_{12}$ and $\eta_{12}\in[\beta_{12},\gamma],\eta_{13}\in[\beta_{13},\gamma],\eta_{23}\in[\beta_{23},\gamma]$. $\qed$
  • Example
    Consider only the antichains $\bot$ and $\{0\}$. Table 3 shows the resulting examples of System 7 and their number of solutions in the column $P_{3}$. The result is Dedekind 3.
  • Example
    Consider only the antichains in [,$[\bot,\{1\}]$. Table 4 shows the resulting examples of System 7 and their number of solutions in the column $P_{3}$. The result is Dedekind 4.
$\alpha$ $\beta_{12}$ $\beta_{13}$ $\beta_{23}$Eq $P_{3}$ $\sum_{\gamma\in[\beta_{12}\vee\beta_{13}\vee\beta_{23},\{0\}]}|[\bot,\alpha]||[\beta_{12},\gamma]||[\beta_{13},\gamma]||[\beta_{23},\gamma]|$ Tot
$\bot$ $\bot$ $\bot$ $\bot$ 1 1 1 + 8 9
$\bot$ $\{0\}$$\{0\}$ $\bot$ 3 1 2 6
$\bot$ $\{0\}$$\{0\}$$\{0\}$ 1 3 1 3
$\{0\}$$\{0\}$$\{0\}$$\{0\}$ 1 1 2 2
$\sum$ 20
Table 3: Simplest example of System 7, the result is D(3)
$\alpha$ $\beta_{12}$ $\beta_{13}$ $\beta_{23}$Eq $P_{3}$ $|[\bot,\alpha]|\times\sum_{\gamma\in[\beta_{12}\vee\beta_{13}\vee\beta_{23},\{1\}]}|[\beta_{12},\gamma]||[\beta_{13},\gamma]||[\beta_{23},\gamma]|$ Tot
$\bot$ $\bot$ $\bot$ $\bot$ 1 1 1 + 8 + 27 = 36 36
$\bot$ $\{0\}$$\{0\}$ $\bot$ 3 1 2 + 12 = 14 42
$\bot$ $\{0\}$$\{0\}$$\{0\}$ 1 3 1 + 8 = 9 27
$\bot$ $\{1\}$$\{1\}$ $\bot$ 3 1 3 9
$\bot$ $\{1\}$$\{1\}$$\{0\}$ 3 2 2 12
$\bot$ $\{1\}$$\{1\}$$\{1\}$ 1 3 1 3
$\{0\}$$\{0\}$$\{0\}$$\{0\}$ 1 1 2 $\times$ (1 + 8) = 18 18
$\{0\}$$\{1\}$$\{1\}$$\{0\}$ 3 1 2$\times$2 = 4 12
$\{0\}$$\{1\}$$\{1\}$$\{1\}$ 1 3 2$\times$1 = 2 6
$\{1\}$$\{1\}$$\{1\}$$\{1\}$ 1 1 3$\times$1 = 2 3
$\sum$ 168
Table 4: One but simplest example of System 9, the result is D(4)
9 A system of eight equations
Let $\alpha_{i},\beta_{i},i\not=j\in\{1,2,3,4\}$ be eight antichains with $\alpha_{i}\leq\beta_{j}$. Consider the set of eight equations in six variables $\chi_{12},\chi_{13},\chi_{14},\chi_{23},\chi_{24},\chi_{34}$, hereafter refered to as System 9:
$\chi_{12}\wedge\chi_{13}\wedge\chi_{14}=\alpha_{1}$ (16)
$\chi_{12}\wedge\chi_{23}\wedge\chi_{24}=\alpha_{2}$ (17)
$\chi_{13}\wedge\chi_{23}\wedge\chi_{34}=\alpha_{3}$ (18)
$\chi_{14}\wedge\chi_{24}\wedge\chi_{34}=\alpha_{4}$ (19)
$\chi_{23}\vee\chi_{24}\vee\chi_{34}=\beta_{1}$ (20)
$\chi_{13}\vee\chi_{14}\vee\chi_{34}=\beta_{2}$ (21)
$\chi_{12}\vee\chi_{14}\vee\chi_{24}=\beta_{3}$ (22)
$\chi_{12}\vee\chi_{13}\vee\chi_{23}=\beta_{4}$ (23)
  • Notation
    We denote by $S_{\alpha}=\alpha_{1}\cup\alpha_{2}\cup\alpha_{3}\cup\alpha_{4},S_{\beta}=\beta_{1}\cup\beta_{2}\cup\beta_{3}\cup\beta_{4},$ the unions of the antichains $\alpha_{i}$ and $\beta_{i}$.
Lemma 9.1
System 9 possibly accepts solutions including sets not in $S_{\alpha}\cup S_{\beta}$.
  • Proof
    We give a counterexample. Let $\beta_{1}=\beta_{4}=\{1234\},\beta_{2}=\beta_{3}=\{123\}$, $\alpha_{1}=\alpha_{2}=\alpha_{3}=\alpha_{4}=\{1\}$ A set of solutions of System 9 is given by $\chi_{12}=\chi_{13}=\chi_{24}=\{1\},\chi_{23}=\{1234\},\chi_{14}=\{123\},\chi_{34}\in[\{1\},\{123\}]$. $\chi_{34}$ can contain sets $3,2$ 3 , 13 , 23 which are not in $S_{\alpha}\cup S_{\beta}$. $\qed$
Due to Lemma 9.1, the number of solutions of System 9 cannot be expected to be computable in polynomial time.
  • Notation
    We denote by $P_{a}(,\alpha_{2},\alpha_{3},\alpha_{4},\beta_{1},\beta_{2},\beta_{3},\beta_{4})$ the number of solutions of System 9 for the given antichains $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ and $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ on the righthand sides of the equations.
10 Application of System 9
System 9 can be used to compute the $(n+4)^{th}$ Dedekind number ($D(n+4)$) based on intervals in the space of antichains over $n$ elements.
Theorem 10.1
Let $n\geq 0$ be a natural number, $N=\{1,..,N\}$, $\mathcal{D}_{n}$ the lattice of antichains over $N$. We have
$D(n+4)=$ $\sum_{\begin{subarray}{c}\alpha_{1}\leq\beta_{2}\wedge\beta_{3}\wedge\beta_{4}\\ \alpha_{2}\leq\beta_{1}\wedge\beta_{3}\wedge\beta_{4}\\ \alpha_{3}\leq\beta_{1}\wedge\beta_{2}\wedge\beta_{4}\\ \alpha_{4}\leq\beta_{1}\wedge\beta_{2}\wedge\beta_{3}\\ \gamma\leq\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3}\wedge\alpha_{4}\\ \delta\geq\beta_{1}\vee\beta_{2}\vee\beta_{3}\vee\beta_{4}\end{subarray}}$
$P_{a}(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\beta_{1},\beta_{2},\beta_{3},\beta_{4})$
$\times$ $|[\gamma,\alpha_{1}]||[\gamma,\alpha_{2}]||[\gamma,\alpha_{3}]||[\gamma,\alpha_{4}]|$
$\times$ $|[\beta_{1},\delta]||[\beta_{2},\delta]||[\beta_{3},\delta]||[\beta_{4},\delta]|$
  • Proof
    We use the decomposition for $\eta\in\mathcal{D}_{n+4}$ in terms of elements of $\mathcal{D}_{n}$:
    $\eta=$ $\eta_{0}\times\{1234\}$
    $\vee\eta_{1}\times\{234\}\vee\eta_{2}\times\{134\}\vee\eta_{3}\times\{124\}\vee\eta_{4}\times{123}$
    $\vee\eta_{12}\times\{34\}\vee\eta_{13}\times\{24\}\vee\eta_{14}\times\{23\}\vee\eta_{23}\times{14}\vee\eta_{24}\times\{13\}\vee\eta_{34}\times{12}$
    $\vee\eta_{123}\times\{4\}\vee\eta_{124}\times\{3\}\vee\eta_{134}\times\{2\}\vee\eta_{234}\times\{1\}$
    $\vee\eta_{1234}\times\{0\}$
    The decomposition is unique iff $A\subseteq B\Rightarrow\eta_{A}\leq\eta_{B}$. We identify $\eta_{12},\eta_{13},\eta_{14},\eta_{23},\eta_{24},\eta_{34}$ with the variables $\chi_{12},\chi_{13},\chi_{14},\chi_{23},\chi_{24},\chi_{34}$ in System 9. Any solution of System 9 for specific values of $\alpha_{i},\beta_{i}$ will now correspond to antichains $\eta$ for which $\eta_{0}=\gamma\leq\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3}\wedge\alpha_{4}$, $\eta_{1234}=\delta\geq\beta_{1}\vee\beta_{2}\vee\beta_{3}\vee\beta_{4}$, $\eta_{i}\in[\gamma,\alpha_{i}],\eta_{jkl}\in[\beta_{i},\delta]$ for $\{i,j,k,l\}=\{1,2,3,4\}$. $\qed$
11 A system of seven equations
Let $\alpha$ and $\beta_{ij}$ with $\alpha\leq\beta_{ij}$ for $\{i,j\}\in\{1,2,3,4\}$ be seven antichains. Note that we will use $ij$ to denote the combination of $i$ and $j$ without any order. We have e.g. $\beta_{ij}\equiv\beta_{ji}$ and $\beta_{1j}\equiv\beta_{j1}$. Consider the set of seven equations in four variables $\chi_{1},\chi_{2},\chi_{3},\chi_{4}$, hereafter referred to as System 11:
$\chi_{1}\wedge\chi_{2}\wedge\chi_{3}\wedge\chi_{4}=\alpha$ (24)
$\chi_{1}\vee\chi_{2}=\beta_{12}$ (25)
$\chi_{1}\vee\chi_{3}=\beta_{13}$ (26)
$\chi_{1}\vee\chi_{4}=\beta_{14}$ (27)
$\chi_{2}\vee\chi_{3}=\beta_{23}$ (28)
$\chi_{2}\vee\chi_{4}=\beta_{24}$ (29)
$\chi_{3}\vee\chi_{4}=\beta_{34}$ (30)
We will use the notation $\beta=\beta_{12}\cup\beta_{13}\cup\beta_{14}\cup\beta_{23}\cup\beta_{24}\cup\beta_{34}$ and $\gamma=\beta-\alpha$. System 11 is a case of System 6 with $r=4$. The number of solutions for System 11 is given by
$P_{4}(\alpha,\beta_{12},\beta_{13},\beta_{14},\beta_{23},\beta_{24},\beta_{34})=\Pi_{\kappa\in C_{\alpha,\gamma}}w(\kappa)$
12 Application of System 11
System 11 can be used to compute the $(n+4)^{th}$ Dedekind number ($D(n+4)$) based on intervals in the space of antichains over $n$ elements.
Theorem 12.1
Let $n\geq 0$ be a natural number, $\mathcal{A}=\{1,..,n\}$, $\mathcal{D}_{n}$ the lattice of antichains over $\mathcal{A}$. We have
$D(n+4)=$ $\sum_{\begin{subarray}{c}\alpha\leq\beta_{12}\wedge\beta_{13}\wedge\beta_{14}\wedge\beta_{23}\wedge\beta_{24}\wedge\beta_{34}\\ \epsilon\geq\delta_{12}\vee\delta_{13}\vee\delta_{14}\vee\delta_{23}\vee\delta_{24}\vee\delta_{34}\\ \beta_{ij}\leq\delta_{ij}\end{subarray}}$
$P_{4}(\alpha,\beta_{12},\beta_{13},\beta_{14},\beta_{23},\beta_{24},\beta_{34})$
$\times$ $P_{4}(\widetilde{\epsilon},\widetilde{\delta_{12}},\widetilde{\delta_{13}},\widetilde{\delta_{14}},\widetilde{\delta_{23}},\widetilde{\delta_{24}},\widetilde{\delta_{34}})$
$\times$ $|[\bot,\alpha_{]}|$
$\times$ $|[\beta_{12},\delta_{12}]|\times|[\beta_{13},\delta_{13}]|\times|[\beta_{14},\delta_{14}]|\times|[\beta_{23},\delta_{23}]|\times|[\beta_{24},\delta_{24}]|\times|[\beta_{34},\delta_{34}]|$
$\times$ |[ϵ,]|italic-$|[\epsilon,\top]|$
  • Proof
    We use the decomposition for $\eta\in\mathcal{D}_{n+4}$ in terms of elements of $\mathcal{D}_{n}$:
    $\eta=$ $\eta_{0}\times\{1234\}$
    $\vee\eta_{1}\times\{234\}\vee\eta_{2}\times\{134\}\vee\eta_{3}\times\{124\}\vee\eta_{4}\times{123}$
    $\vee\eta_{12}\times\{34\}\vee\eta_{13}\times\{24\}\vee\eta_{14}\times\{23\}\vee\eta_{23}\times\{14\}\vee\eta_{24}\times\{13\}\vee\eta_{34}\times\{12\}$
    $\vee\eta_{123}\times\{4\}\vee\eta_{124}\times\{3\}\vee\eta_{134}\times\{2\}\vee\eta_{234}\times\{1\}$
    $\vee\eta_{1234}\times\{0\}$
    The decomposition is unique iff $A\subseteq B\Rightarrow\eta_{A}\leq\eta_{B}$. We identify $\eta_{1},\eta_{2},\eta_{3},\eta_{4}$ with the variables $\chi_{1},\chi_{2},\chi_{3},\chi_{4}$ in System 11, and $\eta_{123},\eta_{124},\eta_{134},\eta_{234}$ with the duals of the variables $\upsilon_{1},\upsilon_{2},\upsilon_{3},\upsilon_{4}$ Any combined solution of the two realisations of System 11 for specific values of $\alpha,\epsilon,\beta_{ij}$ and $\delta_{ij}$ can now be seen to correspond to antichains $\eta$ for which $\eta_{0}\leq\alpha$, $\eta_{1234}\geq\epsilon$, $\beta_{ij}\leq\eta_{ij}\leq\delta_{ij}$. The product of the two $P_{4}$ coefficients counts the number of combined solutions for the $\chi_{i},\upsilon_{i}$$\qed$
Christian Jäkel ([3]) in his computation used a technique based on the decomposition used in the proof of Theorem 12.1. He implemented it on his GPU infrastructure through matrix multiplication. The result in Theorem 12.1 requires to sum over fourteen variables which are constrained by a strong partial order. This reduces the number of terms and allows the sum to be factorised.
  • Example
    We apply Theorem 12.1 to its simplest case of computing $D(4)$. Table 5 shows the non-equivalent values for $\alpha$ and $\beta_{ij}$ withe associated size of the equivalence class and the $P_{4}$ value, which is sometimes zero. Table 6 shows the duals of the corresponding non-equivalent values for $\epsilon$ and $\delta_{ij}$. Finally, table 7 shows the sizes of the combined $\beta-\delta$ equivalence classes. These values allow to compute $D(4)$.
$\alpha$ $\beta_{12}$ $\beta_{13}$ $\beta_{14}$ $\beta_{23}$ $\beta_{24}$ $\beta_{34}$EQ P4
A1 $\bot$ $\bot$ $\bot$ $\bot$ $\bot$ $\bot$ $\bot$ 1 1
[…]
A8 00{0} 00{0} 00{0} 00{0} 00{0} 00{0} 00{0} 1 1
Table 5: Non equivalent values for $\beta_{ij}$ in the system of Theorem 12.1
$\epsilon$ $\delta_{12}$ $\delta_{13}$ $\delta_{14}$ $\delta_{23}$ $\delta_{24}$ $\delta_{34}$EQ P4
B1 00{0} 00{0} 00{0} 00{0} 00{0} 00{0} 00{0} 1 1
[…]
B8 $\bot$ $\bot$ $\bot$ $\bot$ $\bot$ $\bot$ $\bot$ 1 1
Table 6: Non equivalent values for $\widetilde{\delta}_{ij}$ in the system of Theorem 12.1
#EQ B1 B2 B3 B4 B5 B6 B7 B8
A1 1 4 6 1 1
A2
A3
A4 4 4
A5
A6 6
A7 1
A8 1
Table 7: Number of equivalent combinations, leading to a non-zero contribution, in the system of Theorem 12.1
Note the large number of empty cells, corresponding with terms not to be counted in Theorem 12.1. Applying this theorem, the tables 5, 6, 7, leads to the following result, where the exponents in the powers of two correspond to the number of intervals $[\beta_{ij},\delta_{ij}]$ of size two:
$D(4)=2^{6}+2\times(4\times 2^{3}+6\times 2^{1}+4\times 2^{0}+2^{1}+2\times 2^{0})$
$=2^{6}+8\times 2^{3}+14\times 2^{1}+12\times 2^{0}=168$
13 Conclusion
We presented a number of systems of equations in the free distributive lattice of antichains and introduced a technique to count the number of solutions. The systems of equations were chosen to be applicable to the problem of counting the number of antichains of subsets of a finite set of $n$ elements, the $n^{t}h$ Dedekind number $D(n)$ which is the size of the antichain set $\mathcal{D}(n)$. We could apply our technique to reduce $D(n+2)$ to a sum over products of interval sizes in $\mathcal{D}(n)$, which was used in our computation of the ninth Dedekind number. We showed that the technique can be applied for similar reductions of $D(n+3)$ and $D(n+4)$.